A sports trainer wants to know whether the true average time of his athletes who do 100-meter sprint is 98 seconds. He recorder 18 trials of his team and found that the average time is 98.2 seconds with a standard deviation of 0.4 second


1
Expert's answer
2022-06-03T03:53:49-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=98"

"H_1:\mu\not=98"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=17" and the critical value for a two-tailed test is "t_c =2.109816."

The rejection region for this two-tailed test is "R = \{t:|t|>2.109816\}."


The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{98.2-98}{0.4/\sqrt{18}}=2.1213"


Since it is observed that "|t|=2.1213>2.109816=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=17" degrees of freedom, "t=2.1213" is "p=0.048898," and since "p= 0.048898<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 98, at the "\alpha = 0.05" significance level.